daynarussell:

What is the domain of the function y = 2 StartRoot x minus 5 EndRoot? x greater-than-or-equal-to negative 5 x greater-than-or-equal-to 2 x greater-than-or-equal-to 5 x greater-than-or-equal-to 10 @jhonyy9

daynarussell:

[ y = 2\sqrt{x-5} \]

jhonyy9:

\[y = 2\sqrt{(x-5)}\]

daynarussell:

yes

jhonyy9:

ok how you ve learned the term inside radical how not can be ?

daynarussell:

no

jhonyy9:

what no ?

daynarussell:

never mind

jhonyy9:

than you work in a real numbers domain

jhonyy9:

in R

jhonyy9:

when is defined the radical ?

jhonyy9:

please Dayna - this you need learned it

daynarussell:

i have learned it but i forgot sorry

jhonyy9:

can you make radical from a negativ number in real numbers domain ?

jhonyy9:

\[\sqrt{(-1)} = ?\]

jhonyy9:

you ve learned about complex numbers ?

jhonyy9:

in the form of ai +b

jhonyy9:

from where you know that i^2 = -1

jhonyy9:

riri any idea about this ?

jhonyy9:

ok Dayna remember for a radical being defined the term inside radical need being always greater then zero ok ?

jhonyy9:

how you write this condition ?

jhonyy9:

please collaborate

daynarussell:

got it thank you

jhonyy9:

can i check it please ?

jhonyy9:

then no so was my pleasure have a nice day bye bye